OpenStudy (anonymous):

anyone help with Input the equation of the given line in standard form. The line with m = -1/2 and the point (-2, -3).

hartnn (hartnn):

the point slope form of equation of line is \(\large y-y_1=m(x-x_1)\) where m=slope =-1/2 here. point =(x1,y1)= (-2,-3) --->x1= -2 ,y1= -3 just plug in values!

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1}). \] Gradient, \[m=-\frac{ 1 }{ 2 }\] point (-2,-3). \[y-(-3)=(-\frac{ 1 }{ 2 })(x-(-3))\] \[y+3=-\frac{ 1 }{ 2 }x-\frac{ 3 }{ 2 }\] \[y=-\frac{ 1 }{ 2 }x-\frac{ 9 }{ 2 }\]